### v1

# Belief Revision: The Adaptability of Large Language Models Reasoning

Abstract
The capability to reason from text is crucial for real-world NLP applications. Real-world scenarios often involve incomplete or evolving data. In response, individuals update their beliefs and understandings accordingly. However, most existing evaluations assume that language models (LMs) operate with consistent information. We introduce Belief-R111 We will release the dataset and code upon acceptance., a new dataset designed to test LMs’ belief revision ability when presented with new evidence. Inspired by how humans suppress prior inferences, this task assesses LMs within the newly proposed delta reasoning ($\Delta R$) framework. Belief-R features sequences of premises designed to simulate scenarios where additional information could necessitate prior conclusions drawn by LMs. We evaluate $\sim$30 LMs across diverse prompting strategies and found that LMs generally struggle to appropriately revise their beliefs in response to new information. Further, models adept at updating often underperformed in scenarios without necessary updates, highlighting a critical trade-off. These insights underscore the importance of improving LMs’ adaptiveness to changing information, a step toward more reliable AI systems.
Belief Revision: The Adaptability of
Large Language Models Reasoning
Bryan Wiliethanks:  These authors contributed equally., Samuel Cahyawijaya, Etsuko Ishii, Junxian He, Pascale Fung Hong Kong University of Science and Technology Clear Water Bay, Hong Kong bwilie@connect.ust.hk
Figure 1: Belief revision allows reasoners to update their belief based on the new provided evidence. Such ability is necessary to enable better logical reasoning on the case of defeasible inference.
1 Introduction
Human reasoning is characterized by its ability to deal with partial or evolving information. When new information becomes available, we dynamically update our beliefs. We reevaluate and adjust our initial premises or conclusions as necessary in light of this new evidence Łukaszewicz (1990); Brewka (1991). For instance, knowing Tweety is a bird, we conclude that it flies since birds usually fly. Discovering Tweety is a penguin, we retract the conclusion but not the other premises; we still believe Tweety is a bird and that birds typically fly, however, we now conclude that it cannot fly since we know that penguins cannot fly. This form of reasoning permits new information to undermine prior beliefs, which necessitates the ability of belief revision Gärdenfors (1988, 1991); Rott (2001).
Features bAbI 15 FOLIO Proof Writer Leap of Thought $\alpha$NLI BoardgameQA PropInd Belief-R
Incomplete info
Contradictory info
Belief revision
Table 1: The comparison of Belief-R with other widely-used logical reasoning datasets. Belief-R uniquely examines scenarios potentially necessitating belief updates. Belief-R specifically evaluates the capability of belief revision, assessing whether prior beliefs should be adjusted or retained depending of the significance of the new information.
The ability to adjust beliefs allows better adaptability of AI systems by enabling them to properly revise prior inferences as further evidence emerges, such as in commonsense inferences Brewka et al. (1997); Etherington (1986); Pfeifer and Kleiter (2005) and decision-making Antoniou and Williams (1997); Dubois et al. (2002). Despite this, recent reasoning evaluations of state-of-the-art AI technologies, such as language models (LMs), primarily focus on its ability to draw conclusions assuming complete information (c.f. Bhagavatula et al. (2020); Han et al. (2024); Kazemi et al. (2024)). While these evaluations useful to demonstrate the reasoning abilities of LMs, they fail to capture the concept of belief change.
We introduce Belief-R, the first-of-a-kind diagnostic reasoning evaluation dataset designed to assess inferences involving belief revision. Belief-R is inspired by the concept of the Suppression Task Byrne (1989) which enables the retraction of previously inferred beliefs by the introduction of new contextual premises, mimicking how humans reassess their inferences when presented with additional context. To allow a specific and measurable evaluation on belief revision, we introduce a new reasoning evaluation setting dubbed as delta reasoning ($\mathbf{\Delta R}$) framework. Within $\mathbf{\Delta R}$, evaluation is done within two sequential reasoning steps. We start by presenting LMs with two initial premises that satisfy basic logical inference rule to assess its basic inference ability. We expect the model to make accurate inferences to establish the prior beliefs. Then, we introduce another premise to see if the model adjusts its beliefs or keeps them unchanged, depending on the significance of the newly introduced information to the initial beliefs.
Belief-R is specifically designed to support the belief revision evaluation through the $\mathbf{\Delta R}$ framework. Each sample in Belief-R is equipped with two initial premises that support basic modus ponens or modus tollens inferences, and a new premise that brings in new information that might modify previously held beliefs. We synthetically generate the premises in Belief-R leveraging on publicly available dataset, and manually annotate the new information significances along with the ground truth answers through multiple human annotators and majority voting. As illustrated in Figure 2, Belief-R uniquely facilitates thorough evaluations of belief revision capabilities.
Through Belief-R, we evaluate the belief revision ability of small and large scale LMs using different prompting techniques. Our study shows that these models often fail to adjust their responses when presented with new information that necessitates adjustments. We further reveal a critical limitation: they confront a performance trade-off between updating and maintaining their prior beliefs. Models that perform better in the cases where an update is needed, typically faltered on the other. Furthermore, better prompting methods also fail to significantly enhance this capability. These insights underscore a need for strategies to enhance model’s capability to correctly update or maintain its initial beliefs when faced with new evidence to ensure its reliability across evolving scenarios.
2 Related Works
Belief revision
Belief revision is the process of changing beliefs to take into account a new piece of information. In AI systems, one of its early implementation is through procedures by which databases can be updated, i.e. for recording and maintaining reasons for system beliefs Doyle (1979); Falappa et al. (2002); Hansson (2022). Notably, Alchourrón et al. (1985) created formal frameworks to determine how beliefs should be updated in a rational manner. The core challenge in belief revision is deciding rationally which prior beliefs to modify, retain, or discard when confronted with new evidence Rott (2001). Consequently in this paper, we look at how LMs handle belief revision. Belief in LMs can be thought of as models’ output  Li et al. (2019); Jang et al. (2022); Wang et al. (2023a). Several works revise LMs’ beliefs through updating its parameter directly or via fine-tuning De Cao et al. (2021); Dai et al. (2021); Hase et al. (2023). However, this process is not a rational process of the model itself Hofweber et al. (2024). Moreover, it relies on pre-prepared knowledge, which is not ideal if we envision LMs to help with discovering new things Ban et al. (2023); Ma et al. (2024). In this work, we assess LMs’ belief revision capabilities through its response towards queries that neccessitate judgement on whether it needs to update its prior beliefs or keep it.
Language model reasoning evaluation
Reasoning is one of the fundamental intelligent behaviors, essential for solving complex real-world tasks Huang and Chang (2023). Many studies test this behaviour in LMs by setting up various tasks. For instance, some create simple tasks to check if a system can answer questions by connecting facts or using basic logic Weston et al. (2016). Others design more advanced tests to evaluate inductive, deductive, and abductive reasoning Sinha et al. (2019); Saparov et al. (2024); Bhagavatula et al. (2020). Some benchmarks replicate real-world complexities by presenting partial or conflicting informations Talmor et al. (2020); Arabshahi et al. (2021); Sprague et al. (2022); Han et al. (2024); Kazemi et al. (2024). Our research pushes this boundary further, focusing on scenarios where information evolves, presenting queries that neccessitate a dynamic update of prior beliefs in light of this new evidence. We note the comparison in Table 1.
3 Belief Revision
Belief revision is the ability to adapt the reasoning process in response to new information. This capability is critical as it ensures rational decision-making in the face of incomplete and evolving nature of available information Nute (2001); Makinson and Gärdenfors (2005); Ribeiro et al. (2019). In this section, we introduce the concept of belief revision and its notation, and propose the evaluation framework for belief revision capabilities.
3.1 Background and notation
For set of query sentences $\chi$, it encompasses a set of premises $\Gamma{=}\{\gamma_{1},\dots,\gamma_{N}\}$ that could imply a set of conclusions $\Phi{=}\{\varphi_{1},\dots,\varphi_{M}\}$. We denote reasoner’s belief set as a set of sentences $\mathcal{B}$ to represent a contextually fixed background knowledge of $\chi$. In this regard, $\mathcal{B}$ is a tuple that contains set of premises and conclusions: $\mathcal{B}{=}(\Gamma,\Phi)$. In presence of new information $\gamma_{N\texttt{+}1}$, the belief revision concept allow us to infer conclusion $\varphi_{M\texttt{+}1}$ if it is rational to believe $\varphi_{M\texttt{+}1}$ after acknowledging $\gamma_{N\texttt{+}1}$.
Belief revision operation
The belief revision operation is to update belief set $\mathcal{B}$ with a new piece of information, $\gamma_{N\texttt{+}1}$. Here, the result of operation must always be that the beliefs does not contradict one another to avoid inconsistencies among them. The significance of the new information $\gamma_{N\texttt{+}1}$, decides whether it fits with or modifies the existing beliefs after performing the belief revision operation. The operation should smoothly incorporate $\gamma_{N\texttt{+}1}$ and yield a new conclusion $\varphi_{M\texttt{+}1}$ as long as it does not conflict, thereby justifying the maintenance of the reasoner’s prior beliefs. However, if it conflicts, we update the initial beliefs $\mathcal{B}$ appropriately, i.e., by retracting any prior conclusions in $\Phi$, to incorporate the new, conflicting information $\gamma_{N\texttt{+}1}$ to resolve any inconsistencies as we yield the correct $\varphi_{M\texttt{+}1}$. The process to figure out what follows from the revised beliefs is then essentially to infer the new conclusion $\varphi_{M\texttt{+}1}$.
If she has an essay to finish then she will study late in the library
She has an essay to finish
If the library stays open then she will study late in the library
What necessarily had to follow assuming that the above premises were true?
(a) She will study late in the library.
(b) She will not study late in the library.
(c) She may or may not study late in the library. ✓
If she has an essay to finish then she will study late in the library
She has an essay to finish
If she has some textbooks to read then she will study late in the library
What necessarily had to follow assuming that the above premises were true?
(a) She will study late in the library. ✓
(b) She will not study late in the library.
(c) She may or may not study late in the library.
Figure 2: Human reasoning adapts based on new information, leading us to adjust our prior beliefs. Here, the additional condition (left) casts doubt on prior modus ponens conclusion in (a). People may consider that certain other conditions necessary for this conclusion to hold, i.e., the library must remain open. In contrast, the alternative argument (right) does not affect the modus ponens inference pathway, thus prior conclusion could still hold.
3.2 Evaluating belief revision with $\mathbf{\Delta R}$
We introduce a novel delta reasoning ($\mathbf{\Delta R}$) framework, to study how LMs adapt their reasoning when presented with new information over successive timesteps. In this framework, we focus on understanding how model responds to query changes at two essential, consecutive reasoning steps at $t$ and $t\texttt{+}1$. We do this by comparing responses to prior queries at step $t$, $\chi_{t}$, and the next query at step $t\texttt{+}1$, $\chi_{t\texttt{+}1}$, adding the new information $\gamma_{N\texttt{+}1}$.
To begin with, we need $\chi_{t}$ to minimally include two premises, i.e. $\{\gamma_{1},\gamma_{2}\}$, and at least imply conclusion $\varphi_{1}$. We set $\chi_{t}$ to be basic as we expect LMs to answer it in high accuracy to help establish the prior belief and not be affected by the inconsistencies in LMs’ behaviour Jang et al. (2022); Kassner et al. (2021); Hase et al. (2023). We then add the new information $\gamma_{N\texttt{+}1}$ as another premise $\gamma_{3}$ in $\chi_{t\texttt{+}1}$ such that $\chi_{t\texttt{+}1}{=}\{\gamma_{1},\gamma_{2},\gamma_{3}\}$. We examine the corresponding conclusion, $\varphi_{M\texttt{+}1}$, to see how the beliefs shifts according to the significance of $\gamma_{3}$.
One way to set $\chi_{t}$ as basic, is to state them as premises that could satisfy basic logical inference rules of modus ponens and modus tollens Wason and Johnson-Laird (1972); Haack (1978); Evans (1982). Modus ponens and modus tollens is a valid form of inference that have been made a central principle in many propositional and modern logics Copi (1972); Haack (1978). Modus ponens rule of inference states that the premises “if $p$ then $q$” is true and $p$ is true ($p\rightarrow q,p$) satisfy modus ponens conclusion that $q$ must be true ($q$). Modus tollens rule of inference states that the premises “if $p$ then $q$” is true and $q$ is false ($p\rightarrow q,\neg q$) satisfy modus tollens conclusion that $p$ must be false ($\neg p$).
In this setup, we are able to evaluate how well the models revise its beliefs after the introduction of new information in $\gamma_{3}$. We measure the model’s dynamic reasoning ability: whether it can correctly update or maintain its initial beliefs when confronted with new information that may contradict prior beliefs. Through this approach, we can assess both how accurate and how flexible different reasoning models are in evolving scenarios.
Example
Figure 2 presents a scenario where the initial two premises at step $t$, $\gamma_{1}$ and $\gamma_{2}$, adhere to a basic inference rule, modus ponens ($p\rightarrow q,p\vdash q$), implying a $\varphi_{1}$ conclusion of $q$: She will study late in the library. These premises: $\gamma_{1}$, $\gamma_{2}$, and $\varphi_{1}$, form the belief set $\mathcal{B}$. Subsequently, we introduce the third premise $\gamma_{3}$, i.e., another conditional ($r\rightarrow q$) “if the library stays open then she will study late in the library”, as the new information in query $\chi_{t\texttt{+}1}$ and evaluate model’s answer at step $t\texttt{+}1$. This sets the stage to execute the belief revision operation.
Recall $\mathcal{B}$ = $\gamma_{1}$ : If she has an essay to finish then she will study late in the library., $\gamma_{2}$ : She has an essay to finish., $\varphi_{1}$ : She will study late in the library.}, and $\gamma_{3}$ = {If the library stays open then she will study late in the library.}. The introduction of $\gamma_{3}$ suggests that “the library being open” is a sufficient condition for her to “study late in the library”. However, people might consider it as a necessary condition for $\varphi_{1}$. This would involve commonsense reasoning step to recognize that despite the conditions set by $\gamma_{1}$ and $\gamma_{2}$, the actual feasibility of her studying late as concluded in $\varphi_{1}$ might inherently depend on the library’s availability. Thus, while $\gamma_{3}$ does not explicitly redefine the dependency of $\varphi_{1}$ on the library’s status, it implies a scenario where such a dependency could be reasonably inferred. Consequently, we retract $\varphi_{1}$ and infer the new conclusion $\varphi_{2}$: “She may or may not study late in the library”.
4 The Belief-R Dataset
Belief-R is designed to specifically assess the belief revision capability through the $\mathbf{\Delta R}$ framework. To account for this, we adopt a reasoning task that has been extensively studied in cognitive science: the suppression task Byrne (1989). Typically, this task employs a trio of premises $\gamma_{1},\gamma_{2},\gamma_{3}$ that accompanied by three possible conclusions, i.e. as exemplified in Figure 2 for modus ponens: (a) She will study late in the library ($q$), (b) She will not study late in the library ($\neg q$), and (c) She may or may not study late in the library ($\Diamond q\land\Diamond\neg q$; here the symbol $\Diamond$ expresses possibility, $\Diamond q$ can be read as “possibly $q$”).
At step $t$, we form a query $\chi_{t}$ using the first two premises, $\gamma_{1}$ and $\gamma_{2}$. These two premises are the premises that respectively satisfy the modus ponens or modus tollens conclusion, ($p\rightarrow q,p$) or ($p\rightarrow q,\neg q$). These logical rules are basic, and we generally expect that most reasoners can apply them accurately. Next, at step $t\texttt{+}1$, to form the query $\chi_{t\texttt{+}1}$, we introduce a third premise $\gamma_{3}$ which is another conditional statement $r\rightarrow q$. The addition of $\gamma_{3}$ brings in new information that might conflict previously held beliefs. The new information in $r$ can be seen either as adding more requirements or providing an alternative pathway, i.e. to reach the same modus ponens conclusion $q$.
For instance, if $\gamma_{3}$ states if the library stays open then she will study late, we now view $r$: the library stays open as another additional requirement on top of $p$. In such cases, just knowing $p$ alone isn’t enough to conclude $q$: we also need $r$ to be true, thus the condition now becomes $p\land r{\rightarrow}q$. In this case, we retract the prior modus ponens conclusion $q$, and infer the new conclusion $\Diamond q\land\Diamond\neg q$. We refer to this subset of dataset as the “Belief Update” (BU) category. However, in another case, $\gamma_{3}$ could instead states if she has textbooks then she will study late. In this case, $r$ stands as a separate alternative inference path that also leads to $q$, thus $p\lor r{\rightarrow}q$. Here, $p$ still directly leads to $q$, and the acknowledgement of $r$ doesn’t affect this pathway, enabling prior conclusion to still hold. We call this subset as the “Belief Maintain” (BM) category.
In Belief-R, the task requires the model to perform multi-step reasoning to manage the relevance of information within $r$ and decide if it needs to update its prior beliefs at step $t$ or not. The model must discern the implicit commonsense and causal links amongst given premises to identify how $p$ and $r$ are related, determining if their interaction is conjunctive ($p\land r$) or disjunctive ($p\lor r$). Based on the relationships, reasoner needs to determine whether to update its initial conclusion $q$ if the new information $r$ imply an additional requirement for its prior beliefs to hold ($p\land r$), or to maintain its prior beliefs if $r$ simply serves as alternatives ($p\lor r$). To quantitatively measure the model’s reasoning accuracy, we provide multiple choices and ask it to pick the most plausible conclusion. For instance, in examples shown in Figure 2, we would expect LMs to choose options (c) and (a) for each scenario, which aligns with the majority choices made in the original study Byrne (1989); Byrne et al. (1999).
4.1 Dataset construction
We leverage ATOMIC Sap et al. (2019), a publicly-available dataset of everyday commonsense reasoning. It contains textual descriptions of inferential if-then knowledge (e.g., “if X pays Y a compliment, then Y will likely return the compliment”). In addition to the textual commonsense descriptions, the dataset also contains detailed annotation on the type of causal dimensions, i.e. the events, causes (i.e., ‘xIntent’), and effects (i.e., ‘xEffect’, ‘oReact’); with “x” and “o” pertain to PersonX and others.
We use ATOMIC as our seed to ensure the gold-standard validity of our dataset. We synthetically generate Belief-R and minimally introduce variance from the LLM by instructing it to be grounded in the context provided by the seed and not to introduce new ones. We mainly utilize GPT-4 series model as the LLM in our data generation pipeline.
4.1.1 Dataset generation process
We prompt LLM to generate the first two premises conditioned on the events, causes (‘xIntent’, ‘xNeed’, ‘xAttr’), and effects (‘xEffect’, ‘xReact’, ‘xWant’, ‘oEffect’, ‘oReact’, ‘oWant’). We exclude the static elements, as we want to focus on the dynamic causal relationships where change or action is involved, following the original task Byrne (1989). For each event, cause, and effect in ATOMIC, we generate the first two premises in both modus ponens, $p\xrightarrow{}q$ and $p$, and modus tollens, $p\xrightarrow{}q$ and $\neg q$. Afterwards, we prompt LLM to generate the third premises. We design separately the prompt for the alternative and additional conditions (corresponding to the BM and BU categories) within the context in the first premise. For the alternative condition, we prompt the model to generate conditions that are not related at all to $p$ for the conclusions $q$ to happen. For the additional condition, we prompt the model to generate conditions strongly relate to $p$ for this conclusions $q$ to surely hold. Following the original task setup, we set the same third statement in both cases with modus ponens and modus tollens inferences.
In our iterations, we discovered that several entities in the ATOMIC dataset are quite abstract, such as “wants to know what he is selling” or “to analyze the thing in question.” To make these clearer for a general audience and to make them less ambiguous for our study, we prompt LLM to generate more specific examples, changing them to “asks about the price of a pen” or “examine the pen.” To provide more clarity on the dataset generation process, we attach the samples of prompt and generation process in Appendix A. Further, to decide the significance of the third premises, whether it serves as alternative or additional condition, we conducted majority voting among multiple human annotators.
Split Basic @t Belief Update Belief Maintain All w/ 3 premises
Inference rule
Modus ponens 956 537 335 872
Modus tollens 956 537 335 872
Effect entities
Mental states 504 276 184 460
Events 1408 798 486 1284
Total 1912 1074 670 1744
Table 2: Statistics of Belief-R dataset.
Figure 3: Evaluation on basic logical inference capabilities in Belief-R on various LLMs sorted by the #parameters. Pre-trained LLMs with $\geq$6B parameters achieves adequate accuracy ($\geq$75%), while instruction-tuned LLMs achieve the same performance on much smaller scale with $\geq$2.7B parameters.
4.1.2 Ground-truth formulation
To further validate the implied commonsense interaction of the third premises, whether it serves as alternative or additional condition, we manually annotate the final conclusions through a crowdsource annotation task at Appenfootnote 2 (see Appendix B). We cater the variability arises from different interpretations from diverse human readers by asking 5 workers to annotate each problem and then take the majority voting out of them to set the agreed options as the ground truths. Upon further inspection, we found some annotations that logically invalid, i.e. answering $\neg q$ in questions with modus ponens inferences or answering $p$ in modus tollens inferences. We view such cases as non modus ponens (or tollens) inferences and specifically treat the annotation similarly with answering c) $\Diamond q\land\Diamond\neg q$.
In Belief-R, both cases of the logical inferences share the same third statement. To streamline our process, we annotate only the modus ponens samples and then extend the insight on the third premises’ significance to the modus tollens cases. For modus tollens cases, if the corresponding modus ponens sample primarily supports conclusion a) $q$, indicating no conflict with initial beliefs, we set the correct answer to b) $\neg p$. Conversely, if on the modus ponens samples the majority vote suggests the answer c) $\Diamond q\land\Diamond\neg q$, implying additional requirement for the inference, we likewise categorize the corresponding modus tollens cases answers to be c) $\Diamond p\land\Diamond\neg p$. This process maintains the consistencies of the impact of the third premise effectively across related inference scenarios.
4.2 Quality check
Context and logical quality checks
Throughout the data construction phase, we assign one expert to review of the logical formations to ensure they follow the intended structure. We also further gauge the quality of the generated data by reviewing 100 randomly chosen samples to confirm on the context and logical consistency. We conducted a human evaluation via Appen [^2] https://appen.com/, with three native English speakers assessing each sample’s quality. They unanimously confirmed that the conditional relationships in the premises were logically sound across all samples, i.e.  that $q$ entails $p$ and $q$ entails $r$ in both of the conditional premises. We also attach the annotation guidelines in Appendix B.
Dataset filtering
To enhance the quality of our dataset for more reliable evaluation, we refined it by focusing on consensus among annotators. For each question, we utilize the answers manually labeled by five independent workers. We retained only the questions with strong majority agreement (at least four out of five annotators concurred). This filtering retained $\sim$65of the original data.
4.3 Statistics of Belief-R
Table 2 shows the composition of our dataset, sized optimally at around 2K entries to balance representation and computational efficiency for LLM inferences. The dataset includes categories such as Basic @t for basic logical inferences at time $t$, and categories like Belief Update, Belief Maintain, and All w/3 premises for the next step queries at time $t\texttt{+}1$. Additionally, the table details categories inherited from the ATOMIC dataset for the causal relationships of If-Event-Then-Event (e.g., “promoted to senior manager”) and If-Event-Then-Mental-State (e.g., “learns something new”).
Figure 4: BREU score evaluation on belief revision capabilities in Belief-R on various models sorted by the BREU score. While larger-scale LLMs tend to achieve higher BREU score, the performance is far lower compared their basic logical inference at $t$, showcasing limited capability of LLMs in performing belief revision.
5 Experiment Settings
Evaluation metrics
The primary goal of our experiments is to investigate whether LMs possess the capability to perform belief revision in their reasoning processes. We report the accuracies in the Belief Update (BU-Acc) and the Belief Maintain (BM-Acc) subsets to indicate LM capabilities in updating and maintaining its beliefs in which it has to do so. We further introduce a novel metric, BREU (Belief Revision Evaluation Understudy), to assess LMs’ belief revision ability, by averaging BU-Acc and BM-Acc equally. The goal of BREU is to gauge whether the model accurately decides when to update or maintain its prior beliefs. We then benchmark publicly-available LMs and design series of experiments through $\mathbf{\Delta R}$ framework. We perform zero-shot-classification on series of smaller to larger scales pre-trained and finetuned LMs, and prompt LLMs generations through API.
Models
We perform zero-shot classification using encoder-only and decoder-only LMs. For encoder-only LMs, we employ entailment-based inference Yin et al. (2019) using NLI-finetuned LMs of RoBERTa Liu et al. (2019), DeBERTa-v3 base Laurer et al. (2024), and DeBERTa-v3 large Laurer et al. (2023). For decoder-only LMs, we follow Brown et al. (2020) using GPT Radford et al. (2019); Black et al. (2021); Wang and Komatsuzaki (2021), Llama Touvron et al. (2023a, b); AI@Meta (2024), and Phi series Gunasekar et al. (2023); Li et al. (2023); Abdin et al. (2024).
We also include larger-scale LLMs with $\geq$35B parameters. We evaluate the belief revision capability of these larger-scale LLMs via completion API through generation-based approach. We employ three zero-shot prompting methods, i.e., direct prompting (DP), triggering the generation of chain-of-thought (CoT) Kojima et al. (2022), or through plan and solve (PS) prompting Wang et al. (2023b). We employ 8 large-scale LLMs, i.e., Llama-3 70B AI@Meta (2024), Mixtral 8x22B Jiang et al. (2024), Command R, Command R+ Cohere (2024), Claude 3 Haiku, Sonnet Anthropic (2024), GPT-3.5 Turbo, and GPT-4 Turbo OpenAI (2023). Here, we follow Yao et al. (2022) and instruct the model to output the exact character of the final answer as a format. We then retrieve the final answer and report accuracy of the final answer as the metric. When the answer does not follow the format instructed before, we treat it as an instruction-following error.
6 Result and Analysis
(a) Modus Ponens / Modus Tollens
(b) Events / Mental States
(c) Prompting Methods
Figure 5: Performance comparisons dissected across various aspects covering distinction on modus ponens and modus tollens, on different effect entities, and on different prompt methods.
Smaller models fail even on basic logical reasoning tasks.
We start by examining the inferences through the first two premises in Belief-R. In Figure 3, we find as the number of parameters in LMs increases, their ability to handle basic logical inference improves. Smaller models, with <<<2B parameters, struggle with these tasks, scoring close to the majority baseline. Models >>>6B parameters do better, surpassing 75% accuracy. Pre-trained LMs with >{>}>6B parameters achieve $\geq 75\%$ accuracy, while instruction-tuned LMs show an emerging ability from 2.7B parameters achieving significantly higher performance with $\geq 90\%$ accuracy.
LLMs are incapable of revising their prior beliefs.
We group our further exploration on these LMs that performed well ($\geq 75\%$ accuracy) in basic logical inference, and evaluate their average performance in Belief Maintain (BM) and Belief Update (BU) subsets. Despite being a strong reasoner on simple logic, all larger-scale LMs under study fail to perform well on these subsets of Belief-R. In evaluation shown in Figure 4, most of the non-API based models perform almost $0\%$ in BU-Acc, indicating their inability on performing belief revision. We observe that all larger-scale both open-source and commercial LLMs perform better on the belief revision tasks, but their performances are still very limited, achieving at most $\sim$50on BREU.
LLMs confront a trade-off between updating and maintaining their prior beliefs.
We discover a trade-off between BU-Acc and BM-Acc: models performing well on one subset typically faltered on the other, especially in models where the BU-Acc is not close to $0\%$ (see Fig 4). This indicates a potential tension between enhancing specific capabilities, as improving one aspect could inadvertently weaken another. An ideal model would excel at belief revision by consistently making the right decision on whether the new information conflicts with prior beliefs or aligns with them. This underlines the importance of developing strategies that refine the ability to revise beliefs accurately, ensuring its reliability across various scenarios.
7 Discussion
Belief revision is harder in a more complex task with modus tollens inferences.
We compare LLMs’ belief revision capabilities in average, through tasks with modus ponens and modus tollens rule as the basic logical inferences at step $t$. As observed in Figure 5(a), LLMs show reduced BREU score in tasks with modus tollens rule. This is expected, as modus tollens is inherently more difficult relative to modus ponens as it require backward directions of reasoning and it involves reasoning with negations Evans (1982, 1993); Girotto et al. (1997). Furthermore, in tasks involving modus tollens inference, we observe a notably higher BU-Acc compared to a much lower BM-Acc. This disparity suggests that executing accurate belief revision becomes more challenging in complex tasks: decisions to update or maintain beliefs are less clear-cut in these scenarios compared to simpler tasks.
Belief update on mental states effect entities is more challenging than events.
We examine LLMs’ belief revision capabilities in average, when dealing with scenarios involving causal relationships on events and mental states effect entities and note them in Figure 5(b). While the BREU score is similar, LLMs demonstrate tendency towards maintaining their beliefs in mental state effects instead of updating them. This may stem from the challenge of recognizing additional requirements implied from the third, mental state-related, premise which is inherently more abstract and less directly observable than a concrete sequential event.
Better prompting methods yield limited gain on belief revision.
We explore how different prompting methods affect belief revision abilities of LLMs on average. Figure 5(c) shows that CoT, which encourages LLMs to elicit reasoning steps, does not significantly enhance belief revision. While this may stem from its vulnerability to missing-step errors Wang et al. (2023b), attempts to correct these errors with the PS prompting offer minimal benefits, improving only by $\sim$1of BREU. This suggests the ability to revise beliefs could still be absent despite elicitation of reasoning steps.
8 Conclusion
The ability to reason and adapt to changing information is crucial for NLP applications in the real world. Most evaluations assume static knowledge environment, which does not prepare models for dynamic real-life scenarios. To address this, Belief-R is introduced as a diagnostic dataset for evaluating belief revision capability in LMs. Through Belief-R and a novel evaluation framework for evaluating reasoning in a dynamically evolving environment, $\mathbf{\Delta R}$, we reveal that current models struggle with updating their beliefs in response to new information, highlighting the need for improved adaptability and reliability. Our work emphasizes the significance of enhancing the capability of AI models to reason with evolving data for real-world readiness.
Limitations
Towards understanding general belief revision capabilities.
Our study on belief revision using the Belief-R dataset via the $\mathbf{\Delta R}$ framework focuses on belief changes driven by logical inferences like modus ponens and modus tollens, which may not fully represent the complexity of real-world belief revision that often includes a broader range of scenarios and subtleties. Moreover, our methodology primarily considers the introduction of new premises as the trigger for belief revision, overlooking how beliefs might change through re-evaluation of existing knowledge or shifts in perspective in the absence of new information (i.e. in Kronemyer and Bystritsky (2014)). Additionally, our approach to simulate future data is constrained by our inability to determine what LMs have previously known and by resource limitations that restrict the training of large-scale models from scratch.
Intersection of reasoning capability and knowledge capacity.
The evaluation of models’ reasoning capabilities is intricately tied to their knowledge capacity, presenting a significant challenge in discerning pure reasoning capability from mere knowledge recall. Current benchmarks often fail to disentangle these aspects, as models with extensive knowledge bases may appear to possess superior reasoning abilities when, in fact, they might be leveraging stored information rather than demonstrating genuine inferential logic. This conflation complicates the assessment of a model’s true reasoning faculties, as performance improvements on reasoning tasks could be attributed to enhanced information retrieval rather than advancements in reasoning algorithms. Similar to observations in other reasoning datasets, we acknowledge the limitation that the improved performance of models tested on Belief-R might not only stem from their ability to revise beliefs but could also be influenced by superior knowledge recall Huang and Chang (2023). Future research could delve deeper into the relationship between these capabilities, specifically focusing on developing evaluation methods that effectively distinguish between them.
Ethics statement
This research explores how well LMs can revise their beliefs when faced with new information, which is crucial for their use in constantly changing real-world situations. We created a reasoning evaluation dataset to test whether LMs can revise their beliefs correctly or if they stick to their initial assumptions. This is important for using LMs in areas where being accurate and up-to-date is vital, like healthcare or legal advice. In example, being able to revise beliefs appropriately could help prevent LMs from repeating outdated or wrong information, making them more reliable and trustworthy. Plus, LMs that can refresh their understanding according to new societal norms can avoid perpetuating biases, contributing to the fair and ethical use of AI. We consider this a promising and significant area for research. We construct the dataset using events, causes, and effects from ATOMIC and the construction template is designed and reviewed manually and attached in this paper. We utilized crowd-sourced annotators who voluntarily participated through the platform Appenfootnote 2, choosing tasks they deemed fairly compensated. The annotators were presented with multiple-choice tasks predefined to avoid bias and protect privacy, ensuring an ethical annotation process.
Appendix
Appendix A Samples of Prompts
To provide more clarity on the dataset generation process, we attach the samples of prompt in Figure A1.
Figure A1: Samples of prompts utilized in each of the Belief-R generation pipeline. Here, we take the Event: PersonX uses PersonX’s ___ to obtain, Cause (from PersonX): to have an advantage, Effect (to PersonX): pleased from ATOMIC, to generate $p$, $q$, and $r$ for us to form queries at step $t$ and $t\texttt{+}1$ in Belief-R and later go through the manual annotaion process.
Appendix B Annotation guidelines
We provide human annotators with specific guidelines and examples, as detailed in Figures A2 and A3 for ground truth and quality check annotations, respectively.
(a) Annotation Guidelines
(b) Example of Annotation Questions
Figure A2: Details on ground truth annotation
(a) Annotation Guidelines
(b) Example of Annotation Questions
Figure A3: Details on quality check annotation
Appendix C Additional analysis
C.1 LLMs logical reasoning ability are not robust in the presence of distractors
Figure A4: LLMs show decreased inference performance when exposed to noise from the new information in alternative condition.
We analyze the performance of LLMs on basic logical inference tasks and compare it to their accuracy on the BM subset, which differs only by including a third premise. We selected 378 queries from the Belief-R dataset where premises overlap between the basic logical inference tasks at time $t$ and the BM subset for a fair comparison, and visualize them in Figure A4. On most of the models, LMs’ accuracy on samples that do not require change of conclusion (BM-Acc) is dropping compared to its basic inference at $t$ performances. This indicates that the logical reasoning ability of these models are not robust in the presence of distractors, exposing a critical problem of these models especially on the challenges in currently adopted retrieval-augmented-generation (RAG) pipeline to manage noisy documents that have question-related content despite lacking substantive information Lewis et al. (2020); Chen et al. (2022); Gao et al. (2023).
C.2 Better prompting methods yield limited gain on belief revision.
Models Method BU-Acc BM-Acc BREU
Llama-3 Instruct (70B) DP 10.99% 92.09% 51.54%
CoT 12.57% 89.40% 50.99%
PS 12.66% 88.21% 50.44%
Mixtral (8x22B) DP 35.38% 36.57% 35.98%
CoT 27.28% 34.93% 31.11%
PS 44.04% 53.13% 48.59%
Command R DP 12.10% 80.45% 46.28%
CoT 11.36% 81.19% 46.28%
PS 19.37% 69.85% 44.61%
Command R+ DP 13.69% 75.67% 44.68%
CoT 14.71% 77.76% 46.24%
PS 13.41% 65.07% 39.24%
Claude-3 Haiku DP 9.40% 88.66% 49.03%
CoT 13.50% 83.73% 48.62%
PS 13.22% 82.99% 48.11%
Claude-3 Sonnet DP 19.65% 82.69% 51.17%
CoT 21.51% 81.19% 51.35%
PS 16.76% 83.73% 50.25%
GPT-3.5 Turbo DP 14.53% 55.22% 34.88%
CoT 20.48% 65.22% 42.85%
PS 17.78% 67.91% 42.85%
GPT-4 Turbo DP 16.76% 86.72% 51.74%
CoT 13.59% 87.76% 50.68%
PS 12.76% 88.66% 50.71%
Table A1: The effectiveness of various prompting techniques varies across LLMs and subset of Belief-R, enhancing performance in some while degrading it in others.
We provide more details on the investigation in the impact of varied prompting techniques on the performance accuracy of several models, as summarized previously in Figure 5(c). In that figure, the data indicates most significant performance improvements in the BU subset, though overall belief revision improvements remain marginal, showing $\sim$1increase in BREU. In examining the performance across models and different prompting methods as shown in Table A1, it is clear that the influence of these methods is not uniform. For instance, the PS prompting method notably boosted accuracy for models like Mixtral 8x22B and Command R by over 10%. Conversely, this same strategy led to performance reductions in models such as Claude-3 Sonnet and GPT-4 Turbo. Similarly, utilizing CoT and PS exhibited mixed outcomes across models. It strengthened robustness in models like GPT-3.5 Turbo and GPT-4 Turbo, as shown by higher BM-Acc scores, while it increased sensitivity to noise in models like Llama-3 Instruct (70B) and Command R, resulting in reduced BM-Acc values.

### v2

@@
 However, most existing evaluations assume that language models (LMs) operate with consistent information.
-We introduce Belief-R111 We will release the dataset and code upon acceptance., a new dataset designed to test LMs’ belief revision ability when presented with new evidence.
+We introduce Belief-R111 The code and dataset are available at https://github.com/HLTCHKUST/belief-revision, a new dataset designed to test LMs’ belief revision ability when presented with new evidence.
 Inspired by how humans suppress prior inferences, this task assesses LMs within the newly proposed delta reasoning ($\Delta R$) framework.
@@
 Through Belief-R, we evaluate the belief revision ability of small and large scale LMs using different prompting techniques.
-Our study shows that these models often fail to adjust their responses when presented with new information that necessitates adjustments.
+Our study shows that these models are incapable of revising their prior beliefs.
 We further reveal a critical limitation: they confront a performance trade-off between updating and maintaining their prior beliefs.
@@
 Reasoning is one of the fundamental intelligent behaviors, essential for solving complex real-world tasks Huang and Chang (2023).
-Many studies test this behaviour in LMs by setting up various tasks.
-For instance, some create simple tasks to check if a system can answer questions by connecting facts or using basic logic Weston et al.
+One works test this behaviour by creating simple tasks to comprehensively check if a system can answer questions by connecting facts or using basic logic Weston et al.
 (2016).
@@
 (2020).
-Some benchmarks replicate real-world complexities by presenting partial or conflicting informations Talmor et al.
-(2020); Arabshahi et al.
+Some benchmarks replicate real-world complexities by presenting partial or conflicting informations Arabshahi et al.
 (2021); Sprague et al.
@@
 (2024).
-Our research pushes this boundary further, focusing on scenarios where information evolves, presenting queries that neccessitate a dynamic update of prior beliefs in light of this new evidence.
-We note the comparison in Table 1.
+Belief revision focuses on the adaptability problem: whether the model properly revises prior beliefs as new information emerges.
+We extend the reasoning evaluation by focusing on scenarios where information evolves, presenting queries that require dynamic updates of prior beliefs in light of new evidence.
+This is distinct from other existing reasoning tasks involving incomplete or contradictory information, i.e.
+in Talmor et al.
+(2020); Bhagavatula et al.
+(2020); Kazemi et al.
+(2024), since they assume a static environment and focus on filling the gap or resolving the contradiction.
+We further note the comparison in Table 1.
 3 Belief Revision
@@
 For set of query sentences $\chi$, it encompasses a set of premises $\Gamma{=}\{\gamma_{1},\dots,\gamma_{N}\}$ that could imply a set of conclusions $\Phi{=}\{\varphi_{1},\dots,\varphi_{M}\}$.
+In this work, we conceptualize "belief" similarly to its usage in dialogue systems, where it represents what the system currently considers true based on the context Feng et al.
+(2023); van Niekerk et al.
+(2020).
 We denote reasoner’s belief set as a set of sentences $\mathcal{B}$ to represent a contextually fixed background knowledge of $\chi$.
@@
 4.1.2 Ground-truth formulation
-To further validate the implied commonsense interaction of the third premises, whether it serves as alternative or additional condition, we manually annotate the final conclusions through a crowdsource annotation task at Appenfootnote 2 (see Appendix B).
+To further validate the implied commonsense interaction of the third premises, whether it serves as alternative or additional condition, we manually annotate the final conclusions through a crowdsource annotation task at Appensection 4.3 (see Appendix B).
 We cater the variability arises from different interpretations from diverse human readers by asking 5 workers to annotate each problem and then take the majority voting out of them to set the agreed options as the ground truths.
@@
 We also further gauge the quality of the generated data by reviewing 100 randomly chosen samples to confirm on the context and logical consistency.
-We conducted a human evaluation via Appen [^2] https://appen.com/, with three native English speakers assessing each sample’s quality.
+We conducted a human evaluation via Appen22footnotemark: 2, with three native English speakers assessing each sample’s quality.
 They unanimously confirmed that the conditional relationships in the premises were logically sound across all samples, i.e.
@@
 Dataset filtering
-To enhance the quality of our dataset for more reliable evaluation, we refined it by focusing on consensus among annotators.
-For each question, we utilize the answers manually labeled by five independent workers.
-We retained only the questions with strong majority agreement (at least four out of five annotators concurred).
-This filtering retained $\sim$65of the original data.
+To enhance the quality of our dataset for more reliable evaluation, we refine it by focusing on consensus among annotators.
+For each question, we utilize answers manually labeled by five independent workers.
+We measured inter-annotator agreement using Gwet’s AC1 Gwet (2008) [^3] https://pypi.org/project/irrCAC/ since it is better at handling high agreement scenarios, with criteria used in Wongpakaran et al.
+(2013) to interpret the level of agreement.
+We initially measured the annotation agreement which yielded a moderate score of 0.573.
+Further, we observe that some beliefs are naturally more subjective than others.
+To make the evaluation dataset more rigorous, we only take the ones with the higher agreement, as commonly practiced in subjective tasks such as sentiment analysis Bobicev and Sokolova (2017).
+We retain only questions with strong majority agreement (at least 4 out of 5 annotators concurred).
+Post-filtering, as we retain $\sim$65of the original data, the score improved to 0.697 and indicated a substantial agreement.
+Figure 4: BREU score evaluation on belief revision capabilities in Belief-R on various models sorted by the BREU score.
+While larger-scale LLMs tend to achieve higher BREU score, the performance is far below their basic logical inferences at $t$ (Acc@t), showcasing limited capability of LLMs in performing belief revision.
 4.3 Statistics of Belief-R
@@
 Additionally, the table details categories inherited from the ATOMIC dataset for the causal relationships of If-Event-Then-Event (e.g., “promoted to senior manager”) and If-Event-Then-Mental-State (e.g., “learns something new”).
-Figure 4: BREU score evaluation on belief revision capabilities in Belief-R on various models sorted by the BREU score.
-While larger-scale LLMs tend to achieve higher BREU score, the performance is far lower compared their basic logical inference at $t$, showcasing limited capability of LLMs in performing belief revision.
+33footnotetext: https://appen.com/
 5 Experiment Settings
@@
 The primary goal of our experiments is to investigate whether LMs possess the capability to perform belief revision in their reasoning processes.
-We report the accuracies in the Belief Update (BU-Acc) and the Belief Maintain (BM-Acc) subsets to indicate LM capabilities in updating and maintaining its beliefs in which it has to do so.
+Concretely, we consider the models’ predictions on the Basic @t category in the evaluation dataset as the models’ initial belief.
+Thus, changes in accuracy on the Belief Update and Belief Maintain categories directly reflects belief revision.
+We report accuracies in the Belief Update (BU-Acc) and the Belief Maintain (BM-Acc) subsets to indicate LMs’ capabilities in updating and maintaining their beliefs when required.
 We further introduce a novel metric, BREU (Belief Revision Evaluation Understudy), to assess LMs’ belief revision ability, by averaging BU-Acc and BM-Acc equally.
@@
 We perform zero-shot-classification on series of smaller to larger scales pre-trained and finetuned LMs, and prompt LLMs generations through API.
+(a) Modus Ponens / Modus Tollens
+(b) Events / Mental States
+(c) Prompting Methods
+Figure 5: Performance comparisons dissected across various aspects covering distinction on modus ponens and modus tollens, on different effect entities, and on different prompt methods.
 Models
@@
 6 Result and Analysis
-(a) Modus Ponens / Modus Tollens
-(b) Events / Mental States
-(c) Prompting Methods
-Figure 5: Performance comparisons dissected across various aspects covering distinction on modus ponens and modus tollens, on different effect entities, and on different prompt methods.
 Smaller models fail even on basic logical reasoning tasks.
@@
 This disparity suggests that executing accurate belief revision becomes more challenging in complex tasks: decisions to update or maintain beliefs are less clear-cut in these scenarios compared to simpler tasks.
-Belief update on mental states effect entities is more challenging than events.
+Belief update on abstract concept is more challenging for LLMs.
 We examine LLMs’ belief revision capabilities in average, when dealing with scenarios involving causal relationships on events and mental states effect entities and note them in Figure 5(b).
@@
 This may stem from the challenge of recognizing additional requirements implied from the third, mental state-related, premise which is inherently more abstract and less directly observable than a concrete sequential event.
-Better prompting methods yield limited gain on belief revision.
+Better prompting method does not help on belief revision.
 We explore how different prompting methods affect belief revision abilities of LLMs on average.
@@
 This suggests the ability to revise beliefs could still be absent despite elicitation of reasoning steps.
+We put a more detailed analysis in Appendix C.2.
 8 Conclusion
@@
 Through Belief-R and a novel evaluation framework for evaluating reasoning in a dynamically evolving environment, $\mathbf{\Delta R}$, we reveal that current models struggle with updating their beliefs in response to new information, highlighting the need for improved adaptability and reliability.
-Our work emphasizes the significance of enhancing the capability of AI models to reason with evolving data for real-world readiness.
+By exposing these limitations, our work underscores the imperative of developing AI models capable of reasoning adeptly with evolving data, thereby propelling them closer to real-world applicability and robustness.
+9 Future Work
+Observing the limited gain offered by the current prompting strategy, we suggest that future developments aim to enhance LMs’ understanding and handling of premise dependencies.
+A particularly promising direction is to integrate LMs with deterministic symbolic solvers to convert problems into logical rules through chain-of-thought reasoning to aim for a deeper comprehension of the intricate dependencies among different premises.
+Future work could aim to overcome current limitations and pave the way for AI systems to be reliable across evolving scenarios.
+Acknowledgements
+We thank Yejin Bang for the insightful discussions.
+This work has been partially funded by the PF20-43679 Hong Kong PhD Fellowship Scheme and the Hong Kong Fellowship Scheme by the Hong Kong Research Grants Council (RGC), and Care-E Project (FS116).
 Limitations
@@
 Our study on belief revision using the Belief-R dataset via the $\mathbf{\Delta R}$ framework focuses on belief changes driven by logical inferences like modus ponens and modus tollens, which may not fully represent the complexity of real-world belief revision that often includes a broader range of scenarios and subtleties.
-Moreover, our methodology primarily considers the introduction of new premises as the trigger for belief revision, overlooking how beliefs might change through re-evaluation of existing knowledge or shifts in perspective in the absence of new information (i.e.
+To add, our methodology primarily considers the introduction of new premises as the trigger for belief revision, overlooking how beliefs might change through re-evaluation of existing knowledge or shifts in perspective in the absence of new information (i.e.
 in Kronemyer and Bystritsky (2014)).
-Additionally, our approach to simulate future data is constrained by our inability to determine what LMs have previously known and by resource limitations that restrict the training of large-scale models from scratch.
+Furthermore, how to define models’ beliefs is a debatable question and there are some works inspecting models’ internal beliefs by probing the hidden states Burns et al.
+; Zou et al.
+(2023).
+n our work, however, we conceptualize "belief" similarly to its usage in dialogue systems, where it represents what the system currently considers true based on the context Feng et al.
+(2023); van Niekerk et al.
+(2020).
+Further exploration on examining the belief through other approaches (e.g., probing hidden states of LLMs) lies outside this paper’s scope and we leave it as future work.
 Intersection of reasoning capability and knowledge capacity.
@@
 We construct the dataset using events, causes, and effects from ATOMIC and the construction template is designed and reviewed manually and attached in this paper.
-We utilized crowd-sourced annotators who voluntarily participated through the platform Appenfootnote 2, choosing tasks they deemed fairly compensated.
+We utilized crowd-sourced annotators who voluntarily participated through the platform Appensection 4.3, choosing tasks they deemed fairly compensated.
 The annotators were presented with multiple-choice tasks predefined to avoid bias and protect privacy, ensuring an ethical annotation process.
@@
 (2023).
-C.2 Better prompting methods yield limited gain on belief revision.
+C.2 Details on prompting methods variations on their gain at belief revision.
 Models Method BU-Acc BM-Acc BREU